OpenStudy (anonymous):

if a 75kg runner can accelerate from standing start to a speed of 10m/s in 3 s, how much power is generated

OpenStudy (anonymous):

can you please or somebody write a soulution over here

OpenStudy (michele_laino):

please, keep in mind that work done by the runner is equals to the kinetic energy accumulated by that runner, so work=Kinetic energy= \[\frac{ 1 }{ 2 }mV ^{2}=\frac{ 1 }{ 2 }75*10^{2}=3750Joule\] by definition, power is work over time, so: power= \[=\frac{ 3750 }{ 10 }=375Watts\] do you agree?

OpenStudy (michele_laino):

oops...sorry: power= \[=\frac{ 3750 }{ 3 }=1250Watts\] do you agree?